Convert to exponential form. For example, the exponential form of "log5 25 = 2" is "52 = 25".
log e 1 = 0

Answer by oscargut(2103)

log5 (25)=2 becuse 5^2=25
loge (0)=1 because e^0=1
in gral logb(a)=c if and only if b^c=a